OpenStudy (anonymous):

The length of a rectangle is 3 times its width. If the length were increased by 2 inches and the width by 1 inch, the new perimeter would be 62 inches. What is the width and length of the rectangle? If w represents the width of the rectangle, then which equation could be used to solve for w? (w + 1)(3w + 2)=62 2(4w + 3) = 62 2(3w + 3) = 62

OpenStudy (mathmate):

1. Given The length of a rectangle is 3 times its width If the length were increased by 2 inches and the width by 1 inch, the new perimeter would be 62 inches. 2. Unknowns L=length, W=width 3. Equation: L=3W Perimeter = 2(L+2+ W+1)=62 Put in L=3W to eliminate L 2(3W+2+W+1)=62 4. Solve to find W 5. Solution : interpret results Steps 4 and 5 are not necessary. From step 3, can you find the right choice?